OpenStudy (anonymous):

Use the formula for the present value of an ordinary annuity or the amortization formula to solve the following problem. PV=$12,000; i=0.02; PMT=$250; n=? n=____ (Round up to the nearest integer.)

OpenStudy (tkhunny):

First, please present the formula fot the present value of an ordinary annuity. Second, substitute the given values. Third, let's figure out why you're having trouble solving for 'n'.

OpenStudy (anonymous):

\[PV=PMT \frac{ 1-(1+i)^{-n} }{ i }\]

OpenStudy (anonymous):

\[12,000=250\frac{ 1-(1+0.02)^{-n} }{ 0.02 }\]

OpenStudy (tkhunny):

Excellent. Fundamentals are good. Now the algebra. Let's solve for 'n'.

OpenStudy (anonymous):

well i solved all that and got to \[-\frac{ \ln 0.04 }{ \ln1.02 }\]

OpenStudy (anonymous):

i put that into my calculator & got 162.5479173. This is were i get stuck 7 keep getting the wrong answer.

OpenStudy (anonymous):

my homework is a math lad so it tells me if i have the right answer & ever time i do this kinda problem it says i got it wrong.

OpenStudy (tkhunny):

Okay, the algebra looks good. What else can go wrong? Maybe the definition of "Ordinary Annuity"? When is the first $250.00 paid for an "Ordinary Annuity"? Immediately or one year later?

OpenStudy (anonymous):

Idk! 1 year later?

OpenStudy (anonymous):

I GOT IT!! its 163!!

OpenStudy (tkhunny):

That's a little annoying. We did use this symbol "=" and that doesn't quite do it. Anyway, you do have to memorize local conventions. If they mean ">=" when they write "=", we'll just have to get used to it, I guess. Good work.

OpenStudy (tkhunny):

That was next on my list of things that could go wrong. ROUNDING!

OpenStudy (anonymous):

lol THX!!